OpenStudy (anonymous):

please help me!!!! (Solving Exponential Equations with Unequal Bases) Solve 11.32x - 1 = 15.7 (2 points) Answer x ≈ 1.12 x ≈ 0.57 x ≈ 0.94 x ≈ 1.07

OpenStudy (anonymous):

@jazy

OpenStudy (phi):

it would help if you typed the question to show what is in the exponent.

OpenStudy (phi):

but the general idea is if you have \[ a^x= b \] take the log (log base 10 or ln because they are on your calculator) of both sides log(a^x) = log(b) x log(a)= log(b) x= log(b)/log(a) that is the general idea. Can you use it to do your problem ?

OpenStudy (anonymous):

yes; I got x ≈ 0.94 @phi

OpenStudy (phi):

exactly what is the exponent in this problem. I was guessing 2x-1 ? if so, I did not get x=0.94

OpenStudy (anonymous):

what did you get? :/ @phi

OpenStudy (phi):

what is the exponent ?

OpenStudy (johnweldon1993):

\[11.3^{2x-1} = 15.7\] @InsanelyChaotic this is what @phi is asking...does the equation look like this?